# Existence of Solutions to the Seiberg-Witten Vortex Equations with Exponential Decay on the Plane

### v1

@@ 3.1. Gauge Theory Set-Up
$\displaystyle\iint_{\mathbf{R}^{2}}d(\phi_{2}(d\phi_{1}-iA\phi_{2})-\phi_{1}(d\phi_{2}+iA\phi_{1}))$
$=\lim_{R\to\infty}\iint_{|x|\leq R}d(\phi_{2}(d\phi_{1}-iA\phi_{2})-\phi_{1}(d\phi_{2}+iA\phi_{1}))$
@@
Therefore, the integral of $d(\phi_{2}(d\phi_{1}-iA\phi_{2})-\phi_{1}(d\phi_{2}+iA\phi_{1}))$ over $\mathbf{R}^{2}$ is equal to zero. Hence, by Lemma 3.2, we immediately get the estimate of $\mathscr{Y}$ as claimed. The statement of equality can be checked directly via calculations (also, see [MR0614447], Ch.3). ∎
@@ 3.2. Seiberg-Witten Gauge Theory
where $A_{j}:\mathbf{R}^{4}\to\mathbf{R}$ and $\psi:\mathbf{R}^{4}\to\mathbf{C}$. The associated connection on $L_{\rho}$ is given by $A=iA_{0}\,dx_{0}+iA_{1}\,dx_{1}+iA_{2}\,dx_{2}+iA_{3}\,dx_{3}$. Note that $\psi:\mathbf{R}^{4}\to\mathbf{C}$ is called a spinor and $\nabla_{A}$ is called a spinor connection. We denote $\mathcal{C}(S^{+})=i\Omega^{1}(\mathbf{R}^{4})\times C^{\infty}(\mathbf{R}^{4},\mathbf{C}^{2})$ by the configuration of the $spin^{c}$ structure $S^{+}$.
@@ 3.3. Dimensional Reduction
$\Rightarrow\begin{bmatrix}\dfrac{\partial}{\partial x_{0}}\psi_{1}+iA_{0}\psi_{1}-i\dfrac{\partial}{\partial x_{1}}\psi_{1}+A_{1}\psi_{1}-iA_{2}\psi_{2}+A_{3}\psi_{2}\\ \dfrac{\partial}{\partial x_{0}}\psi_{2}+iA_{0}\psi_{2}+i\dfrac{\partial}{\partial x_{1}}-A_{1}\psi_{2}+iA_{2}\psi_{1}+A_{3}\psi_{1}\end{bmatrix}=0$
@@
(3.10) $\begin{cases}2\overline{\partial}\,\Phi=-\dfrac{1}{2}\psi_{1}\overline{\psi_{2}}\,dz\wedge d\overline{z}\\ F(A)=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2})\,dz\wedge d\overline{z}\\ \begin{bmatrix}-\dfrac{1}{2}\Phi^{0,1}&-(\overline{\partial}+A^{0,1})\\ \partial+A^{1,0}&-\dfrac{1}{2}\Phi^{1,0}\end{bmatrix}\cdot\begin{bmatrix}\psi_{1}\\ \psi_{2}\end{bmatrix}=0.\end{cases}$
@@ 4.1. A Kazdan-Warner Equation
We would like to comment that (4.4) can only make sense at the points in $\mathbf{C}$ where $\psi_{1}$ and $\psi_{2}$ do not vanish. Denote $S_{n}=\{z_{1},\cdots,z_{n}\}$ by the set of all zeroes of $\psi_{1}$ and $\psi_{2}$. Let $\theta_{j}$ be a multivalued function such that $\psi_{j}=\exp\left(\dfrac{1}{2}|\psi_{j}|^{2}+\dfrac{i}{2}\theta_{j}\right)$. Then $\log(\overline{\psi_{1}}\psi_{2})=\dfrac{1}{2}(|\psi_{1}|^{2}+|\psi_{2}|^{2}-i(\theta_{1}-\theta_{2}))$. For now, assume $\psi_{1},\psi_{2}$ are such that the only solution to (4.4) are constants holomorphic function on $\mathbf{C}$. Then for each constant $C\in\mathbf{C}$, we must have $|\psi_{1}|^{2}+|\psi_{2}|^{2}-i(\theta_{1}-\theta_{2})=C$. If we write $C=C_{1}+iC_{2}$, then $|\psi_{1}|^{2}=C_{1}-|\psi_{2}|^{2}$. Given this, as we substitute the expression for $\alpha$ in (4.3) into the remaining equation of the system (4.2), we obtain
@@
Denote by $Vor_{\mathfrak{e}}(\mathbf{C})$ the moduli space of solutions of (4.1) with property (E). There is well-defined map from $\eta_{\mathfrak{e}}:Vor_{\mathfrak{e}}(\mathbf{C})\to\bigcup_{n\in\mathbf{N}}Sym^{n}(\mathbf{C})$. It is interesting to investigate more about the properties of such map $\eta_{\mathfrak{e}}$. We will address this direction elsewhere.
Remark 4.3.
Note that the construction of the solutions at the end of the previous section is not of property (E). In fact, they all have polynomial growth. Recall that the collection of the gauge equivalence classes of such solutions is denoted by $Vor_{\mathfrak{p}}(\mathbf{C})$. Similarly, there is an obvious map $\eta_{\mathfrak{p}}:Vor_{\mathfrak{p}}(\mathbf{C})\to\bigcup_{n\in\mathbf{N}}Sym^{n}(\mathbf{C})$. It is not difficult to see that $\eta_{\mathfrak{p}}$ is surjective (cf. Theorem 3.11).
The first part of Theorem 4.1 is an immediate consequence of the following theorem.
@@
The proof of Theorem 4.4 will be given in the following subsections. Next, we discuss some aspects that go into the proof. The strategy we employ here is similar to that of Taubes’ in [MR0614447] and will be divided into several steps. (cf. 4.1)
@@ 4.2. Some Estimates for $\mathscr{A}$
By definition, $\left\lVert\nabla v\right\rVert^{2}_{L^{2}}\leq\left\lVert v\right\rVert^{2}_{L^{2}_{1}}$. For the second term, without loss of generality, we suppose $v$ is a function with compact support. Then by the Hölder inequality, there is a positive constant $c(v)>0$ and some large $c>0$ such that
$\left\lVert v\right\rVert_{L^{1}}\leq c(v)\cdot\left\lVert v\right\rVert_{L^{2}}\leq e^{c\left\lVert v\right\rVert^{2}_{L^{2}}}\leq e^{c\left\lVert v\right\rVert^{2}_{L^{2}_{1}}}.$
@@ 4.3. Infinimum of $\mathscr{A}$
The proof is a standard bootstrapping argument and the fact that solutions to elliptic equations are real analytic. ∎

### v2

@@ 3.1. Gauge Theory Set-Up
$\displaystyle\iint_{\mathbf{R}^{2}}d(\phi_{1}(d\phi_{0}-iA\phi_{1})-\phi_{0}(d\phi_{1}+iA\phi_{0}))$
$=\lim_{R\to\infty}\iint_{|x|\leq R}d(\phi_{1}(d\phi_{0}-iA\phi_{1})-\phi_{0}(d\phi_{1}+iA\phi_{0}))$
@@
Therefore, the integral of $d(\phi_{1}(d\phi_{0}-iA\phi_{1})-\phi_{0}(d\phi_{1}+iA\phi_{0}))$ over $\mathbf{R}^{2}$ is equal to zero. Hence, by Lemma 3.2, we immediately get the estimate of $\mathscr{Y}$ as claimed. The statement of equality can be checked directly via calculations (also, see [MR0614447], Ch.3). ∎
@@ 3.2. Seiberg-Witten Gauge Theory
where $A_{j}:\mathbf{R}^{4}\to\mathbf{R}$ and $\psi:\mathbf{R}^{4}\to\mathbf{C}^{2}$. The associated connection on $L_{\rho}$ is given by $A=iA_{0}\,dx_{0}+iA_{1}\,dx_{1}+iA_{2}\,dx_{2}+iA_{3}\,dx_{3}$. Note that $\psi:\mathbf{R}^{4}\to\mathbf{C}^{2}$ is called a spinor and $\nabla_{A}$ is called a spinor connection. We denote $\mathcal{C}(S^{+})=i\Omega^{1}(\mathbf{R}^{4})\times C^{\infty}(\mathbf{R}^{4},\mathbf{C}^{2})$ by the configuration of the $spin^{c}$ structure $S^{+}$.
@@ 3.3. Dimensional Reduction
$\Rightarrow\begin{bmatrix}\dfrac{\partial}{\partial x_{0}}\psi_{1}+iA_{0}\psi_{1}-i\dfrac{\partial}{\partial x_{1}}\psi_{1}+A_{1}\psi_{1}-iA_{2}\psi_{2}+A_{3}\psi_{2}\\ \dfrac{\partial}{\partial x_{0}}\psi_{2}+iA_{0}\psi_{2}+i\dfrac{\partial}{\partial x_{1}}\psi_{2}-A_{1}\psi_{2}+iA_{2}\psi_{1}+A_{3}\psi_{1}\end{bmatrix}=0$
@@
(3.10) $\begin{cases}2\overline{\partial}\,\Phi=-\dfrac{1}{2}\psi_{1}\overline{\psi_{2}}\,dz\wedge d\overline{z}\\ F(A)=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2})\,dz\wedge d\overline{z}\\ \begin{bmatrix}\dfrac{1}{2}\Phi^{0,1}&-(\overline{\partial}+A^{0,1})\\ \partial+A^{1,0}&-\dfrac{1}{2}\Phi^{1,0}\end{bmatrix}\cdot\begin{bmatrix}\psi_{1}\\ \psi_{2}\end{bmatrix}=0.\end{cases}$
@@ 4.1. A Kazdan-Warner Equation
We would like to comment that (4.4) can only make sense at the points in $\mathbf{C}$ where $\psi_{1}$ and $\psi_{2}$ do not vanish. Denote $S_{n}=\{z_{1},\cdots,z_{n}\}$ by the set of all zeroes of $\psi_{1}$ and $\psi_{2}$. Let $\theta_{j}$ be a multivalued function such that $\psi_{j}=\exp\left(\dfrac{1}{2}|\psi_{j}|^{2}+\dfrac{i}{2}\theta_{j}\right)$. Then $\log(\overline{\psi_{1}}\psi_{2})=\dfrac{1}{2}(|\psi_{1}|^{2}+|\psi_{2}|^{2}-i(\theta_{1}-\theta_{2}))$. For now, assume $\psi_{1},\psi_{2}$ are such that the only solution to (4.4) are constant functions on $\mathbf{C}$. Then for each constant $C\in\mathbf{C}$, we must have $|\psi_{1}|^{2}+|\psi_{2}|^{2}-i(\theta_{1}-\theta_{2})=C$. If we write $C=C_{1}+iC_{2}$, then $|\psi_{1}|^{2}=C_{1}-|\psi_{2}|^{2}$. Given this, as we substitute the expression for $\alpha$ in (4.3) into the remaining equation of the system (4.2), we obtain
@@
Denote by $Vor_{\mathfrak{e}}(\mathbf{C})$ the moduli space of solutions of (4.1) with property (E). There is a well-defined map from $\eta_{\mathfrak{e}}:Vor_{\mathfrak{e}}(\mathbf{C})\to\bigcup_{n\in\mathbf{N}}Sym^{n}(\mathbf{C})$, where each gauge equivalence class of a solution $(A_{0},A_{1},\psi_{1},\psi_{2})$ of (4.1) is mapped to an unordered set of zeroes of $\psi_{1}$ and $\psi_{2}$, which we view as an element of some symmetric product of $\mathbf{C}$. It is interesting to investigate more about the properties of such map $\eta_{\mathfrak{e}}$. We will address this direction elsewhere.
Theorem 4.1 is an immediate consequence of the following theorem.
@@
The proof of Theorem 4.3 will be given in the following subsections. Next, we discuss some aspects that go into the proof. The strategy we employ here is similar to that of Taubes’ in [MR0614447] and will be divided into several steps (cf. 4.1). The statement of Theorem $4.4$ might be well-known to experts. However, we could not find a direct reference to it in the literature. Thus, we include a detailed proof of it here for our sake and the sake of readers.
@@ 4.2. Some Estimates for $\mathscr{A}$
By definition, $\left\lVert\nabla v\right\rVert^{2}_{L^{2}}\leq\left\lVert v\right\rVert^{2}_{L^{2}_{1}}$. For the second term, without loss of generality, we suppose $v$ is a function with compact support. Then by the Hölder inequality, there is a positive constant $b>0$ that depends only on $v$ and some large constant $c>0$ such that
$\left\lVert v\right\rVert_{L^{1}}\leq b\cdot\left\lVert v\right\rVert_{L^{2}}\leq e^{c\left\lVert v\right\rVert^{2}_{L^{2}}}\leq e^{c\left\lVert v\right\rVert^{2}_{L^{2}_{1}}}.$
@@ 4.3. Infinimum of $\mathscr{A}$
The proof is a standard bootstrapping argument and the fact that solutions to elliptic equations are real analytic, see, e.g., Theorem 4.5 in Section 3.4 of [MR0614447] or Morrey’s book [MR202511] for more details. ∎